# Bridging Relevance and Reasoning: Rationale Distillation in Retrieval-Augmented Generation

### v1

@@ Abstract
The reranker and generator are two critical components in the Retrieval-Augmented Generation (i.e., RAG) pipeline, responsible for ranking relevant documents and generating responses. However, due to differences in pre-training data and objectives, there is an inevitable gap between the documents ranked as relevant by the reranker and those required by the generator to support answering the query. To address this gap, we propose RADIO, a novel and practical preference alignment framework with RAtionale DIstillatiOn. Specifically, We first propose a rationale extraction method that leverages the reasoning capabilities of Large Language Models (LLMs) to extract the rationales necessary for answering the query. Subsequently, a rationale-based alignment process is designed to rerank the documents based on the extracted rationales, and fine-tune the reranker to align the preferences. We conduct extensive experiments on two tasks across three datasets to demonstrate the effectiveness of our approach compared to baseline methods. Our code is released online  [^1] The link will be released after publication to ease reproduction.
@@
Pengyue Jia1111equal contribution, Derong Xu1111equal contribution, Xiaopeng Li1111equal contribution, Zhaocheng Du2, Xiangyang Li2, Xiangyu Zhao1, Yichao Wang2, Yuhao Wang1, Huifeng Guo2, Ruiming Tang2, 1City University of Hong Kong, 2Huawei Noah’s Ark Lab
@@ 1 Introduction
Large Language Models (LLMs), pretrained on massive datasets, have demonstrated exceptional reasoning and text generation capabilities, as evidenced by prior research Zhao et al. (2023). These models also adhere to the scaling laws, exhibiting improvements in performance and intelligence as the number of parameters increases Kaplan et al. (2020). Retrieval-augmented generation (RAG) builds upon these capabilities by integrating information retrieval mechanisms with generative models, such as LLMs. This approach not only mitigates the problem of hallucination in text generation but also enhances the system’s adaptability to dynamically evolving information needs, making it a robust solution for tasks requiring both accuracy and contextual relevance Gao et al. (2023).
@@
However, the RAG pipeline typically assembles components (e.g., the reranker and generator Fan et al. (2024)) that have been pretrained separately. Due to differences in their pretraining data and optimization objectives, these components often exhibit varying preferences within the pipeline, which can impact the overall effectiveness of the system. Specifically, pretrained rerankers Xiao et al. (2023) excel at evaluating the relevance between queries and documents. However, the documents identified as “relevant” under this criterion may not provide the necessary support for reasoning to derive an accurate answer to the query. Bridging this gap between the reranker’s relevance-based rankings and the generator’s requirements for reasoning support presents a significant challenge that must be addressed to improve the RAG pipeline’s performance. Recent studies try to address this gap by training a bridge model Ke et al. (2024) or using LLM-based Zhang et al. (2024) and retrieval-based scores to fine-tune RAG components Dong et al. (2024). Additionally, while some other methods are not specifically designed to tackle this problem, they can indirectly help bridge the gap. These approaches can use response quality Ma et al. (2023) or perplexity distillation Izacard et al. (2023); Shi et al. (2023) as signals to fine-tune the reranker. While these approaches show promise, they face critical limitations: relying solely on the connection between the query or answer and the document to determine the document’s importance captures only surface-level relationships, neglecting the deeper, rationale-based reasoning.
@@
To this end, we propose RADIO, a novel and practical preference alignment framework with rationale distillation in retrieval-augmented generation. First, to efficiently extract the rationales necessary for answering a query, we leverage the query and its ground truth answer as contextual information. Then, we generate the rationale required to derive the correct answer with LLMs. Second, to mitigate the preference misalignment between the reranker and generator while ensuring the solution remains practical, we rerank the documents based on the extracted rationale and fine-tune the reranker using this rationale-based ranking. This step distillates rationales from generators to rerankers, and aligns the reranker with the generator’s information needs for answering the query effectively.
@@
RADIO effectively addresses the preference inconsistency between RAG components by first generating a comprehensive rationale and then fine-tuning the reranker based on the extracted rationale. This approach considers the deeper reasoning behind answers. We evaluate RADIO on two tasks across three datasets: Open-domain QA (NQ Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017)) and Multi-choice questions (MMLU Hendrycks et al. (2020)). The results validate the superiority of our method compared to other state-of-the-art baselines. Our contributions can be summarized as follows:
@@
  • Extensive experiments are conducted on two tasks across three datasets to demonstrate the effectiveness and transferability of RADIO.
@@ 2 Methodology
In this section, we introduce rationale distillation from generators to rerankers to mitigate the preference misalignment between rerankers and generators in RAG. Specifically, we first demonstrate the task definition of RAG in Section 2.1. Then we give an overview of our proposed framework in Section 2.2, introduce the rationale extraction method in Section 2.3, and detail the rationale-based alignment in Section 2.4 and optimization in Section 2.5.
@@ 2.1 Task Definition
To address the hallucination problem and enhance adaptability to dynamic information of LLMs, RAG systems have been proposed. These systems enhance generative models by introducing additional contextual information retrieved based on a given query $q$. Specifically, when a query $q$ is input into the RAG pipeline, the retriever $\mathcal{R}_{\text{retriever}}$ first identifies relevant documents by calculating similarity scores and top-$k_{1}$ selection. The process can be formalized as follows:
@@ 2.2 Framework Overview
In the rationale extraction process, we aim to obtain the correct rationale required to answer the query accurately. To achieve this, we combine the query with the ground truth answer and input them into LLM. By using the correct answer as guiding context in the prompt, we can obtain the rationale with greater precision. This approach leverages the LLM’s reasoning capabilities more effectively, ensuring that the generated rationale closely aligns with the requirements for deriving the correct answer.
In the rationale-based alignment process, our goal is to use the generated rationale to guide the reranker, enabling it to select documents that better support the generator in answering the query. Specifically, we leverage the generated rationale as a signal to rerank documents. The reranked documents will be used to fine-tune the reranker, addressing the preference misalignment between the reranker and the generator. By aligning these components, the process ensures that the selected documents are not only contextually relevant but also optimally supportive for the generator’s reasoning and response generation.
@@ 2.3 Rationale Extraction
Rationales are critical components of LLM reasoning processes and have been shown to significantly enhance the accuracy of LLM-generated responses. This perspective is supported by existing works such as Chain-of-Thoughts (CoT Wei et al. (2022)) and O1 Zhong et al. (2024). Existing work Shi et al. (2023); Ma et al. (2023); Dong et al. (2024) has primarily focused on the initial relationships between queries and documents or indirect relationships between answers and documents, while overlooking rationales—a crucial intermediary component in the reasoning process. Motivated by this, we aim to extract rationales as signals to align the preferences of different components in the RAG pipeline.
To accurately extract the rationale necessary for answering a query and deriving the correct answer, we combine the query with the ground truth answer as contextual information of LLMs, as shown in Figure 1(a). The prompt template used for this process is as follows: "You are a professional QA assistant. Given a question and the ground truth answer, you can output the rationale why the ground truth answer is correct. Question: {question}. Answer: {answer}. Rationale: ". By doing so, we effectively bridge the gap between the query and the answer by generating the necessary rationale. This rationale accurately supports the reasoning process required to derive the correct answer from the query.
@@ 3.1 Datasets and Metrics
To evaluate RADIO with other methods, we conduct experiments on two tasks across three datasets: Open-domain QA (NQ Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017)) and Multi-choice questions (MMLU Hendrycks et al. (2020)). Deatiled dataset descriptions are given in Appendix A.1. Following previous work Ma et al. (2023); Shi et al. (2023), we report EM and F1 scores for Open-domain QA datasets and EM for MMLU.
@@ 3.2 Baselines
Table 1: Overall experiments.
@@
DPA-RAG 0.3391 0.4674 0.3385 0.4710 0.3462 0.4793 0.6080 0.7076 0.6097 0.7119 0.6149 0.7169
RADIO (Ours) 0.3512 0.4790 0.3565 0.4850 0.3665 0.4917 0.6084 0.7095 0.6128 0.7137 0.6154 0.7151
Method MMLU (Source Dataset NQ) MMLU (Source Dataset TriviaQA)
Humanities Social STEM Other ALL Humanities Social STEM Other ALL
RADIO (Ours) 0.4172 0.7013 0.5080 0.6717 0.5562 0.423 0.6942 0.5090 0.6678 0.5559
@@ 3.3 Backbone Rerankers
To validate the generality and adaptability of RADIO, we selected three different rerankers as the backbone models for our experiments: (1) gte-base Li et al. (2023): a reranker model proposed by Alibaba DAMO Academy, with 109M parameters and 768 embedding dimensions. (2) gte-large Li et al. (2023): the larger version of gte-base, with 335M parameters and 1024 embedding dimensions. (3) bge-reranker-base Xiao et al. (2023): a powerful cross-encoder architecture reranker proposed by Beijing Academy of Artificial Intelligence, with 278M parameters.
@@ 3.4 Implementation Details
We implement RADIO on FlashRAG Jin et al. (2024), a Python library for efficient RAG research. In the RAG pipeline, we take e5-base-v2 Wang et al. (2022) as the retriever, and Meta-Llama-3.1-8B-Instruct Touvron et al. (2023) as the generator. We sample 20,000 instances from NQ and TriviaQA to construct the fine-tuning dataset and fine-tune rerankers separately. For document sampling, we set the shift $n$ in Top-k shifted by N method as 3, and sample 6 negative samples from the subsequent documents. To ensure a fair comparison, the sampling index is fixed and remains unchanged across methods. In the RAG pipeline, we set the number of documents selected by retriever and reranker (i.e., $k_{1}$ and $k_{2}$) as 20 and 5. For fine-tuning the reranker, we tune the training epochs from 1 to 5 and the integration hyperparameter $\alpha$ from 0.0 to 1.0. We use Adam Kingma (2014) optimizer with a learning rate 6e-5 and a weight decay of 0.01. The prompts we used in experiments are given in Appendix A.4.
@@ 3.5.1 Open-domain QA
To evaluate the effectiveness of RADIO and its transferability across different rerankers, we conduct experiments on the NQ and TriviaQA datasets. The results are presented in Table 1. From these results, we can draw the following conclusions:
@@
  • Compared to other baseline methods, RADIO consistently achieves superior performance across all datasets and reranker backbone configurations. The results validate the effectiveness of using rationales as signals for preference alignment.
  • On TriviaQA, methods such as RRR and REPLUG show performance declines relative to the base method when using rerankers gte-large and bge-reranker-base. This indicates that these methods are sensitive and have limited applicability. In contrast, RADIO demonstrates robust adaptability to different rerankers, achieving significant performance improvements across all three rerankers.
@@
3.5.2 MMLU
We also conduct experiments on MMLU. Since MMLU is a multiple-choice dataset, we report the EM metric Ma et al. (2023). Additionally, following previous work Yu et al. (2023), we fine-tune the reranker using open-domain QA as the source task and evaluate its performance on the MMLU dataset. The results are shown in Table 2 and we can draw the following conclusions:
  • From the metrics corresponding to the ALL category, RADIO demonstrates consistent improvements over the base method. This highlights the effectiveness and transferability of RADIO, as it successfully adapts to multi-choice question tasks even when fine-tuned on the Open-domain QA tasks.
  • Analyzing the results by question category, RADIO shows more significant improvements over the base method in the Humanities and Social Sciences categories, with average gains of 2.74% and 1.61%, respectively. However, it exhibits a slight negative effect in the STEM category. This may be due to the fine-tuning datasets (NQ and TriviaQA), which are Open-domain QA datasets with distributions more similar to humanities and social sciences but markedly different from STEM subjects.

### v2

@@ Abstract
The reranker and generator are two critical components in the Retrieval-Augmented Generation (i.e., RAG) pipeline, responsible for ranking relevant documents and generating responses. However, due to differences in pretraining data and objectives, there is an inevitable misalignment between the documents ranked as relevant by the reranker and those required by the generator to support query-specific answers. To bridge this gap, we propose RADIO, a novel and practical preference alignment framework with RAtionale DIstillatiOn. Specifically, we first propose a rationale extraction method that leverages the reasoning capabilities of Large Language Models (LLMs) to extract the rationales necessary for answering a query. Subsequently, a rationale-based alignment process is designed to rerank documents based on the extracted rationales and fine-tune the reranker to better align the preferences. Extensive experiments conducted on three tasks across four datasets demonstrate the effectiveness and transferability of our approach. Our code is released online [^1] https://github.com/Applied-Machine-Learning-Lab/RADIO.
@@
Pengyue Jia1111Equal contribution, Derong Xu1,2111Equal contribution, Xiaopeng Li1111Equal contribution, Zhaocheng Du3, Xiangyang Li3 Yichao Wang3, Yuhao Wang1, Qidong Liu1, Maolin Wang1, Huifeng Guo3 Ruiming Tang3222Corresponding author, Xiangyu Zhao1222Corresponding author 1City University of Hong Kong, 2University of Science and Technology of China, 3Huawei Noah’s Ark Lab Correspondence: xianzhao@cityu.edu.hk
@@ 1 Introduction
Large Language Models (LLMs), pretrained on massive datasets, have demonstrated exceptional reasoning and text generation capabilities, as evidenced by prior research Zhao et al. (2023); Ni et al. (2025); Wang et al. (2023b); Xu et al. (2024); Fu et al. (2023). These models also adhere to the scaling laws, exhibiting improvements in performance and intelligence as the number of parameters increases Kaplan et al. (2020). Retrieval-augmented generation (RAG) Xu et al. (2025a, b) builds upon these capabilities by integrating information retrieval mechanisms Zhao et al. (2018a, b); Liu et al. (2024) with generative models, such as LLMs. This approach not only mitigates the problem of hallucination in text generation but also enhances the system’s adaptability to dynamically evolving information needs, making it a robust solution for tasks requiring both accuracy and contextual relevance Gao et al. (2023).
@@
However, RAG pipelines typically assemble components (e.g., the reranker and generator Fan et al. (2024)) that have been pretrained separately. Due to differences in their pretraining data and optimization objectives, these components often exhibit varying preferences, which can impact the overall effectiveness of the system. Specifically, pretrained rerankers Xiao et al. (2023) are designed to evaluate the relevance between queries and documents. However, the documents identified as relevant under this criterion may not provide the necessary support for reasoning to derive an accurate answer to the query. Bridging this gap between the reranker’s relevance measurement and the generator’s reasoning requirements presents a significant challenge that must be addressed to improve the RAG pipeline’s performance.
Recent studies try to address this gap by training a bridge model Ke et al. (2024), using LLM-based scores Zhang et al. (2024) or combining both LLM-based and retrieval-based scores Dong et al. (2024) to fine-tune RAG components. Additionally, while some other methods are not explicitly designed for this problem, they can indirectly contribute to bridging the gap. These approaches can use response quality Ma et al. (2023) or perplexity distillation Izacard et al. (2023); Shi et al. (2023) as signals to fine-tune the reranker. Despite showing promise, these approaches face critical limitations: their alignment signals rely solely on the surface-level connection between the query/answer and document, failing to capture the deeper reasoning processes or more complex relationships involved.
@@
To address the above limitation, we propose RADIO, a novel and practical preference alignment framework with rationale distillation in RAG. RADIO leverages rationale as a signal to bridge the reranker’s relevance measurement with the generator’s reasoning requirements for response generation. First, to efficiently extract the rationales needed to answer a query, we use the query and its ground truth answer as context and generate the rationales with LLMs. Second, to mitigate the preference misalignment between the reranker and generator while ensuring the solution remains practical, we rerank the documents based on the extracted rationales and fine-tune the reranker. This step distillates rationales from generators to rerankers, and aligns the reranker with the generator’s information needs for answering the query effectively.
Figure 1: Overview of RADIO.
@@
RADIO effectively addresses the preference inconsistency between RAG components by first generating a comprehensive rationale and then fine-tuning the reranker based on the extracted rationale. This approach considers the deeper reasoning behind answers. We evaluate RADIO on three tasks across four datasets: Open-domain QA (NQ Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017)), Multi-choice questions (MMLU Hendrycks et al. (2020)), and Multi-hop QA (Musique Trivedi et al. (2022)). The results validate the superiority of our method compared to other state-of-the-art baselines. Our contributions can be summarized as follows:
@@
  • Extensive experiments are conducted on three tasks across four datasets to demonstrate the effectiveness and transferability of RADIO.
2 Related Work
Retrieval-Augmented Generation. Large language models (LLMs) have demonstrated groundbreaking performance across numerous tasks Jia et al. (2025) but still face challenges such as hallucination and outdated knowledge Gao et al. (2023). To address these issues, retrieval-augmented generation (RAG) has been introduced Fan et al. (2024); Zhang et al. (2025); Jia et al. (2024a). RAG retrieves relevant information from external knowledge bases and incorporates it as contextual input to the generator (LLM), enhancing the accuracy and reliability of the generated responses. The typical RAG pipeline can be divided into several key components: query rewriter, retriever, reranker, and generator. The query rewriter Wang et al. (2023a) modifies and expands Jia et al. (2024b); Li et al. (2023a) the original query to improve retrieval recall, ensuring more relevant documents are retrieved. The retriever Chen et al. (2024) fetches relevant documents based on the query. Dense retrievers generally outperform sparse retrievers in this step. To integrate contextual information more effectively and identify documents more relevant to the query, rerankers Moreira et al. (2024a) with larger models and greater complexity are introduced to reorder the retrieved documents compared to retrievers. Finally, the generator—usually a powerful LLM such as GPT-4 Achiam et al. (2023) or Llama Touvron et al. (2023)—uses the query and the top-$k$ documents from the reranker to generate the final response. In this work, we address the issue of preference misalignment among different components within the RAG pipeline. We aim to leverage rationale as a signal to align these preferences and enhance the overall performance of the RAG system. It is worth noting that RADIO and methods like ReAct Yao et al. (2022) are not mutually exclusive and can be used together to complement each other. While RADIO enhances the ranking of retrieved documents at each step, ReAct focuses on task-solving through reasoning and action.
Preference Alignment. To further improve LLMs, preference alignment is often performed after the initial pretraining phase Jiang et al. (2024). Approaches such RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2024) are proposed to align the output of LLMs more closely with human preferences. DPO transforms tasks into classification problems, achieving high computational efficiency and strong performance. In the context of RAG, several works Ke et al. (2024) can be transformed to address the challenge of preference alignment between RAG components. REPLUG Shi et al. (2023) improves RAG pipelines involving black-box LLMs by using the probability of the LLM generating the correct answer as a signal to determine document importance. Similarly, RRR Ma et al. (2023) uses metrics based on the quality of the LLM’s generated response as a signal to evaluate a document’s utility. On the other hand, ARL2 Zhang et al. (2024) prompts LLMs to generate self-guided relevance labels for fine-tuning retriever, and DPA-RAG Dong et al. (2024) introduces a bidirectional alignment strategy to mitigate preference inconsistencies in RAG pipelines. In this work, we focus on optimizing the reranker within RAG. Our goal is to enable the reranker to effectively identify supportive documents well-suited for the generator, whether black-box or open-source, facilitating the production of accurate outputs. In addition, BGM Ke et al. (2024) trains a bridge model between retriever and LLMs to transform the retrieved information into the format LLM’s prefer. Our method is theoretically compatible with BGM and could be combined to jointly enhance performance without any conflict.
@@ 3 Methodology
In this section, we detail rationale distillation in RAG. Specifically, we first demonstrate the task definition of RAG in Section 3.1. Then we give an overview of our proposed framework in Section 3.2, introduce the rationale extraction method in Section 3.3, and detail the rationale-based alignment in Section 3.4 and optimization in Section 3.5.
@@ 3.1 Task Definition
To address the hallucination problem and enhance adaptability to dynamic information of LLMs, RAG systems have been proposed. These systems enhance generative models by introducing additional contextual information retrieved based on a given query $q$. Specifically, when a query $q$ is input into the RAG pipeline, the retriever $\mathcal{R}_{\text{retriever}}$ first retrieves relevant documents by calculating similarity scores and top-$k_{1}$ selection. The process can be formalized as follows:
@@ 3.2 Framework Overview
In the rationale extraction process, we combine the query with its ground truth answer and input them into LLMs to generate precise rationales. Using the correct answer as context in the prompt improves the accuracy of the LLM’s rationale generation, ensuring that the generated rationale closely aligns with the requirements for deriving the correct answer.
In the rationale-based alignment process, our goal is to use the generated rationale to guide rerankers, enabling it to select documents that better support the generator in answering the query. Specifically, we leverage the generated rationale as a signal to rerank documents. The reranked documents will be used to fine-tune the reranker, addressing the preference misalignment between the reranker and the generator. By aligning these components, the process ensures that the selected documents are not only contextually relevant but also optimally supportive for the generator’s reasoning and response generation.
@@ 3.3 Rationale Extraction
Rationales are critical components of LLM reasoning processes and have been shown to enhance the accuracy of LLM-generated responses significantly. This perspective is supported by existing works such as Chain-of-Thoughts (CoT Wei et al. (2022)). Existing work Shi et al. (2023); Ma et al. (2023); Dong et al. (2024) has primarily focused on the initial relationships between queries and documents or indirect relationships between answers and documents, while overlooking rationales, a crucial intermediary component in the reasoning process. Motivated by this, we aim to extract rationales as signals to align the preferences of different components in the RAG pipeline.
To accurately extract the rationale necessary for answering a query and deriving the correct answer, we combine the query with the ground truth answer as contextual information of LLMs, as shown in Figure 1(a). The prompt template used for this process is as follows: "You are a professional QA assistant. Given a question and the ground truth answer, you can output the rationale why the ground truth answer is correct. Question: {question}. Answer: {answer}. Rationale: ". The generation process can be formalized as:
$r=\text{LLM}(q,a)$ (4)
where $q$ and $a$ are the query and answer, $r$ deontes the generated rationale.
By doing so, we effectively bridge the gap between the query and the answer by generating the necessary rationale. This rationale accurately supports the reasoning process required to derive the correct answer from the query.
@@ 4.1 Datasets and Metrics
Following previous work Shi et al. (2023); Zhang et al. (2024), we conduct experiments on three tasks across four datasets to evaluate RADIO with other methods: Open-domain QA (NQ Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017)), Multi-choice questions (MMLU Hendrycks et al. (2020)), and Multi-hop QA (Musique Trivedi et al. (2022)). Detailed dataset descriptions are given in Appendix A.1. Following previous work Ma et al. (2023); Shi et al. (2023), we report EM and F1 scores for QA datasets and EM for MMLU.
@@ 4.2 Baselines
Table 1: Overall experiments. “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. For all metrics, higher is better. $\Delta$ represents the relative improvement of RADIO over Base method.
@@
DPA-RAG 0.3391 0.4674 0.3385 0.4710 0.3462 0.4793 0.6080 0.7076 0.6097 0.7119 0.6149 0.7169*
RADIO (Ours) 0.3512* 0.4790* 0.3565* 0.4850* 0.3665* 0.4917* 0.6084 0.7095* 0.6128* 0.7137* 0.6154 0.7151
$\Delta$ $\uparrow 19.82$% $\uparrow 18.39$% $\uparrow$27.41$\uparrow$23.25$\uparrow$8.72$\uparrow$6.82$\uparrow$11.55$\uparrow$11.31$\uparrow$11.52$\uparrow$11.27$\uparrow$0.65$\uparrow$0.40%
Table 2: Experimental results on MMLU. EM is reported as the evaluation metric. The source dataset used to fine-tune rerankers is the Open-domain QA dataset NQ. $\Delta$ represents the relative improvement of RADIO over Base method.
Method Humanities Social STEM Other ALL
Base 0.4089 0.6867 0.5147 0.6650 0.5502
Atlas 0.3985 0.6935 0.5074 0.6563 0.5447
REPLUG 0.4102 0.6854 0.5065 0.6590 0.5473
RRR 0.4079 0.6913 0.5116 0.6572 0.5484
ARL2 0.4147 0.7016 0.5106 0.6630 0.5540
DPA-RAG 0.4157 0.701 0.5078 0.6652 0.5541
RADIO (Ours) 0.4172 0.7013 0.5080 0.6717 0.5562
$\Delta$uparrow2.03$\uparrow$2.13$\downarrow$1.30$\uparrow$1.01$\uparrow$1.09%
@@ 4.3 Backbone Rerankers
To validate the generality and adaptability of RADIO, we select three different rerankers as the backbone models for our experiments: gte-base Li et al. (2023b), gte-large Li et al. (2023b), and bge-reranker-base Xiao et al. (2023). A detailed introduction is given in Appendix A.3.
@@ 4.4 Implementation Details
We implement RADIO on FlashRAG Jin et al. (2024), a Python library for efficient RAG research. In the RAG pipeline, we take e5-base-v2 Wang et al. (2022) as the retriever, and Meta-Llama-3.1-8B-Instruct Touvron et al. (2023) as the generator. We sample 20,000 instances from NQ and TriviaQA to construct the fine-tuning dataset and fine-tune rerankers separately. For document sampling, we set the shift $n$ in Top-k shifted by N method as 3, and sample 6 negative samples from the subsequent documents. To ensure a fair comparison, the sampling index is fixed and remains unchanged across methods. In the RAG pipeline, we set the number of documents selected by retriever and reranker (i.e., $k_{1}$ and $k_{2}$) as 20 and 5. The retrieval corpus we used in experiments is Wikipedia (2018, December). For fine-tuning the reranker, we tune the training epochs from 1 to 5 and the integration hyperparameter $\alpha$ from 0.0 to 1.0. We use Adam Kingma (2014) optimizer with a learning rate 6e-5 and a weight decay of 0.01. The prompts we used in experiments are given in Appendix A.4. We provide an analysis of the impact of fine-tuning sample size on performance in Appendix A.5, and we also give a detailed analysis on the shift parameter $n$ and the negative samples $m$ in Appendix A.6.
@@ 4.5.1 Open-domain QA
To evaluate the effectiveness of RADIO and its transferability across different rerankers, we conduct experiments on the NQ and TriviaQA datasets. The results are presented in Table 1. A case study is given in Appendix A.7. From these results, we can draw the following conclusions:
@@
  • Compared to other baseline methods, RADIO consistently achieves superior performance across all datasets and reranker backbone configurations. The results validate the effectiveness of using rationales as signals for preference alignment in RAG pipeline.
  • On TriviaQA, methods such as RRR and REPLUG show performance declines relative to the base method when using rerankers gte-large and bge-reranker-base. This indicates that these methods are sensitive, limiting their applicability. In contrast, RADIO demonstrates robust adaptability to different rerankers, achieving significant performance improvements across all three rerankers.
@@
4.6 Task Generalization Analysis
4.6.1 Multi-choice Questions
We also conduct experiments on MMLU. Since MMLU is a multiple-choice dataset, we report the EM metric Ma et al. (2023). Additionally, following previous work Yu et al. (2023), we fine-tune the reranker bge-reranker-base using open-domain QA as the source task and evaluate its performance on the MMLU dataset. Table 2 shows the results of fine-tuning reranker with NQ dataset. The results of fine-tuning reranker with TriviaQA are given in Appendix A.8. We can draw the following conclusions:
  • From the metrics corresponding to the ALL category, RADIO demonstrates consistent improvements over Base. This highlights the effectiveness and generalization of RADIO, as it successfully adapts to multi-choice question tasks even when fine-tuned on the Open-domain QA tasks.
  • Analyzing the results by question category, RADIO shows more significant improvements over the Base method in the Humanities and Social Sciences categories, with average gains of 2.03% and 2.13%, respectively. However, it exhibits a slight negative effect in the STEM category. This may be due to the fine-tuning datasets (NQ and TriviaQA), which are Open-domain QA datasets with distributions more similar to humanities and social sciences but markedly different from STEM subjects.